# Efficient4D: Fast Dynamic 3D Object Generation from a Single-view Video

### v3

@@ Abstract
Generating dynamic 3D object from a single-view video is challenging due to the lack of 4D labeled data. An intuitive approach is to extend previous image-to-3D pipelines by transferring off-the-shelf image generation models such as score distillation sampling. However, this approach would be slow and expensive to scale due to the need for back-propagating the information-limited supervision signals through a large pretrained model. To address this, we propose an efficient video-to-4D object generation framework called Efficient4D. It generates high-quality spacetime-consistent images under different camera views, and then uses them as labeled data to directly reconstruct the 4D content through a 4D Gaussian splatting model. Importantly, our method can achieve real-time rendering under continuous camera trajectories. To enable robust reconstruction under sparse views, we introduce inconsistency-aware confidence-weighted loss design, along with a lightly weighted score distillation loss. Extensive experiments on both synthetic and real videos show that Efficient4D offers a remarkable 10-fold increase in speed when compared to prior art alternatives while preserving the quality of novel view synthesis. For example, Efficient4D takes only 10 minutes to model a dynamic object, vs 120 minutes by the previous art model Consistent4D. Our code is publicly available at https://github.com/fudan-zvg/Efficient4D.
@@ 3 Preliminary
For rendering, each Gaussian includes opacity $\alpha$ and view-dependent color $\mathbf{c}$ represented by spherical harmonics (SH). Given an arbitrary view $\mathcal{I}$ defined by intrinsic and extrinsic parameters, we render the pixel at position $(u,v)$ with timestamp $t$ by blending visible Gaussians:
@@
with $\text{with }T_{i}=$ $\prod^{i-1}_{j=1}(1-f_{j}(u,v,t)\alpha_{j}).$ (3)
@@ 4.1 Image synthesis across views and timestamps
where $i\in\{1,2\}$ and $w$ denotes the fusion ratio, we systematically vary the ratio $w$ from 0 to 0.5, resulting in distinct columns of “Feature fusion" as illustrated in Figure 3. For $w=0$, the original generation process is represented, where the spatial volumes of the two frames are independent, leading to temporal inconsistency. As $w$ approaches $0.5$, we observe a gradual convergence of the generated astronauts conditioned on different input images, achieving similarity in both texture and geometry. This suggests that smoothing at the feature level is effective in aligning frames over time.
However, a challenge arises as the fused features may induce similar motion. While the bottom-left astronaut is stepping forward, the one at the bottom-right column does not exhibit forward motion like the top astronaut. Thus, a trade-off is necessary between achieving temporal consistency and preserving motion independence. In practical terms, it is recommended to set the ratio just above $0.5$, striking a balance that ensures temporal texture consistency at a moderate cost of entangled geometry. Also note that the results in Figure 3 are not sensitive to the choice.
@@ 5.1 Experiment setup
Implementation details In equation (8), the spatial volumes are smoothed locally in a sliding window style, so the consistency of generated images may be weakened as the frame number increase. In practice, we can adjust the weight following two rules: (i) the middle weight should be at least $0.5$; (ii) At least half of the volumes should have a positive weight. Here we provide some sample weights which have a good effect:
@@
where $N$ denotes the frame number. The weight will be normalized to ensure the sum is $1$.
@@
In the reconstruction stage, Gaussians are initialized randomly inside the sphere with radius space 0.5 with identity rotations and a initial number of 5,000. Training of 4D Gaussian Splatting is carried out using the Adam optimizer for 1500 iterations with batch size 1. All other 4Dgs hyperparameters remain consistent with those in [15]. The balancing weights in loss function (equation 13) are set as $\tilde{\lambda}_{\text{rgb}}=8000,\tilde{\lambda}_{\text{ssim}}=2000,\lambda_{\text{SDS}}=1$. In equation (12), the calculation of estimated image $\mathbf{I}$ is implemented by frame interpolation using a optical flow based model [48]. For each generated frame $I_{t}$, we use its adjacent four frames to interpolate the estimated frame as follows:
@@ 5.5 Sparse input case
We assessed our method’s performance in an extremely sparse input scenario, comprising only two discrete frames. In such cases, we set $w=0.25$ in equation (9). We also modified the code of Consistent4D [14] to make it adapt to such case. As illustrated in Figure 7, our approach successfully generates images featuring smooth motion and high spatiotemporal consistency. In contrast, Consistent4D fails to operate effectively under such conditions.
Consider a scenario where we seek 4D modeling for a static toy. While the toy can take different poses, it lacks autonomous movement, posing a challenge for continuous video capturing. In these instances, our method demonstrates effectiveness by requiring only a few key frames to produce dynamic content, thereby expanding the potential applications of the 4D generation task.
@@ 5.6.1 Ablation of image generation
Figure 8: Ablation study on image generation, time-synchronous spatial volumes and frame interpolation. The images follow a chronological order from left to right.
@@
We first performed ablation studies to assess the influence of different components in our image generation stage. To better evaluate the image quality itself, we reconstruct the 4D Gaussian without SDS loss. We compared our full method against three baseline settings: (1) Only input video are used for reconstruction; (2) Time-synchronous spatial volume is excluded; (3) Frame interpolation is excluded. Note that setting (2) is equivalent to the case where SyncDreamer-T is replaced with original SyncDreamer. The results corresponding to these settings can be found in Figure 8. Table 3 also gives quantitative evaluations.
@@ 5.6.2 Ablation of reconstruction stage
In the reconstruction stage, we will study the effect of three components: (1) confidence map, (2) supervision of generated images, and (3) SDS loss. The visual results are shown in Figure 9 and 10.

### v4

@@ Abstract
Generating dynamic 3D object from a single-view video is challenging due to the lack of 4D labeled data. An intuitive approach is to extend previous image-to-3D pipelines by transferring off-the-shelf image generation models such as score distillation sampling. However, this approach would be slow and expensive to scale due to the need for back-propagating the information-limited supervision signals through a large pretrained model. To address this, we propose an efficient video-to-4D object generation framework called Efficient4D. It generates high-quality spacetime-consistent images under different camera views, and then uses them as labeled data to directly reconstruct the 4D content through a 4D Gaussian splatting model. Importantly, our method can achieve real-time rendering under continuous camera trajectories. To enable robust reconstruction under sparse views, we introduce inconsistency-aware confidence-weighted loss design, along with a lightly weighted score distillation loss. Extensive experiments on both synthetic and real videos show that Efficient4D offers a remarkable 10-fold increase in speed when compared to prior art alternatives while preserving the quality of novel view synthesis. For example, Efficient4D takes only 10 minutes to model a dynamic object, vs 120 minutes by the previous art model Consistent4D. Our code and video results are publicly available at https://fudan-zvg.github.io/Efficient4D/.
@@ 2 Related work
Recent works have attempted to generate 4D videos from a single input video. For example, TrajectoryCrafter [46] leverages rendered point cloud as the condition of diffusion model to facilitate novel view synthesis. However, the point cloud is projected from the source video and its estimated depth, so there are no points from the back view, limiting the range of novel views. Another work ReCamMaster [47] trains a diffusion model which introduces frame dimension conditioning for input video and camera pose conditioning for target trajectory. But the trajectory is still limited to the front 180. In contrast, our goal is to generate the complete 360 views for a dynamic object.
@@ 3 Preliminary
[…]
@@
$\text{with }T_{i}=$ $\prod^{i-1}_{j=1}(1-f_{j}(u,v,t)\alpha_{j}).$ (3)
@@ 4.1 Image synthesis across views and timestamps
where $i\in\{1,2\}$ and $w$ denotes the fusion ratio, we systematically vary the ratio $w$ from 0 to 0.5, resulting in distinct columns of “Feature fusion” as illustrated in Figure 3. For $w=0$, the original generation process is represented, where the spatial volumes of the two frames are independent, leading to temporal inconsistency. As $w$ approaches 0.50.50.5, we observe a gradual convergence of the generated astronauts conditioned on different input images, achieving similarity in both texture and geometry. This suggests that smoothing at the feature level is effective in aligning frames over time.
However, a challenge arises as the fused features may induce similar motion. While the bottom-left astronaut is stepping forward, the one at the bottom-right column does not exhibit forward motion like the top astronaut. Thus, a trade-off is necessary between achieving temporal consistency and preserving motion independence. In practical terms, it is recommended to set the ratio just above 0.50.50.5, striking a balance that ensures temporal texture consistency at a moderate cost of entangled geometry. Also note that the results in Figure 3 are not sensitive to the choice.
@@ 5.1 Experiment setup
Implementation details In equation (8), the spatial volumes are smoothed locally in a sliding window style, so the consistency of generated images may be weakened as the frame number increase. In practice, we can adjust the weight following two rules: (i) the middle weight should be at least 0.50.50.5; (ii) At least half of the volumes should have a positive weight. Here we provide some sample weights which have a good effect:
@@
where $N$ denotes the frame number. The weight will be normalized to ensure the sum is 111.
@@
In the reconstruction stage, Gaussians are initialized randomly inside the sphere with radius space 0.5 with identity rotations and a initial number of 50,000 without densification. Training of 4D Gaussian Splatting is carried out using the Adam optimizer for 500 iterations with batch size 1. All other 4Dgs hyperparameters remain consistent with those in [15]. The balancing weights in loss function (equation 13) are set as $\tilde{\lambda}_{\text{rgb}}=8000,\tilde{\lambda}_{\text{ssim}}=2000,\lambda_{\text{SDS}}=1$. In equation (12), the calculation of estimated image $\mathbf{I}$ is implemented by frame interpolation using a optical flow based model [50]. For each generated frame $I_{t}$, we use its adjacent four frames to interpolate the estimated frame as follows:
@@ 5.5 Challenging cases
Figure 8: Performance on cases including rotation. We show two novel views for each case.
We assessed our method’s performance on two challenging cases: extremely sparse input and rotation dynamics. In sparse case, the input comprising only two discrete frames. In such cases, we set $w=0.25$ in equation (9). We also modified the code of Consistent4D [14] to make it adapt to such case. As illustrated in Figure 7, our approach successfully generates images featuring smooth motion and high spatiotemporal consistency. In contrast, Consistent4D fails to operate effectively under such conditions. Consider a scenario where we seek 4D modeling for a static toy. While the toy can take different poses, it lacks autonomous movement, posing a challenge for continuous video capturing. In these instances, our method demonstrates effectiveness by requiring only a few key frames to produce dynamic content, thereby expanding the potential applications of the 4D generation task.
For another rotation case, we adopt the data introduced in [61]. As shown in the first row of Figure 8, the character will rotate its body in the monocular video input, which is complex and challenging. However, our approach can also successfully infer the accurate geometry while Consistent4D tends to generate fragments. For example, from the Consistent4D results in Figure 8, we observe that the punching figure and the kicking man both suffer fragmented arms or legs.
@@ 5.6.1 Ablation of image generation
Figure 9: Ablation study on image generation, time-synchronous spatial volumes and frame interpolation. For each case, we show the input images in the first row and three novel-view images generated by SyncDreamer-T at different timestamps in the other rows, followed by one rendered image from the reconstructed 4DGS at the middle timestamp. Each column refers to a specific timestep. For the row “No interp”, there is no interpolated input image in the middle, thus leading to a vacancy at timestamp $t_{2}$.
@@
Figure 10: Ablation study on the temporal synchronization. We compare SV3D [62], SyncDreamer [8], SyncDreamer with consistent self-attention [63] (SyncDreamer-A) and our proposed Syncdreamer-T. We show nine consecutive frames for each case to better evalute the temporal consistency.
We first performed ablation studies to assess the influence of different components in our image generation stage. To better evaluate the image quality itself, we reconstruct the 4D Gaussian without SDS loss. We compared our full method against three baseline settings: (1) Only input video are used for reconstruction; (2) Time-synchronous spatial volume is excluded; (3) Frame interpolation is excluded. Note that setting (2) is equivalent to the case where SyncDreamer-T is replaced with original SyncDreamer. The results corresponding to these settings can be found in Figure 9. Table 3 also gives quantitative evaluations. Beyond the above components, we also ablate the design of temporal synchronization in Figure 10.
@@
[…]
Ablation on temporal synchronization We also examine the design of temporal synchronization. In conceptual terms, directly applying image-to-3D methods to each frame independently is inferior due to intrinsic generative randomness. This can be validated in Figure 10 by comparing SV3D [62], SyncDreamer [8] and SyncDremaer-T(Ours). The image-to-3D methods (SV3D and SyncDreamer) cannot generate a consistent shape or texture across frames. Therefore, a proper temporal synchronization mechanism is necessary. In Section 4.1, we propose SyncDreamer-T, which blends the 3D volume feature. We also implement a consistent self-attention mechanism introduced in [63] in SyncDreamer [8], resulting in Syncdreamer-A. Specifically, in each self-attention block that projects $Q_{i},K_{i},V_{i}$ corresponding to the image $I_{i}$, we randomly sample another image $I_{j}$’s $K_{j},V_{j}$, and concatenate two $K$s and $V$s as the new $K_{i}$ and $V_{i}$. The results in Figure 10 indicate that our method generates a more consistent texture and smoother motion than SyncDreamer-A. We ascribe this to the following reason: The cross-attended features may focus too much on the details that its control on global texture and motion is weakened.
@@ 5.6.2 Ablation of reconstruction stage
Figure 13: Ablation study on the choice of 4D representation. We apply the same optimization strategy to three 4D representations: Cascade-DyNeRF [14], Deformable-3DGS [44] and 4DGS [15]. 4DGS features the fastest convergence speed.
In the reconstruction stage, we will study the effect of three components: (1) confidence map, (2) supervision of generated images, (3) SDS loss, and (4) choice of 4D representation. The visual results are shown in Figure 11, 12 and 13.
@@
[…]
Choice of 4D representation Our synthetic training data is versatile and supports the optimization of 4D representation models, such as Cascade-DyNeRF [14] and Deformable-3DGS [44] used in other 4D generation works [14, 53]. In the comparison shown in Figure 13, 4DGS [15] converge in 500 iterations while the others do not. Also, thanks to the fast convergence speed of 4DGS, our method achieves significant efficiency advantage.